The monthly precipitation (in inches) for May is listed for 10 different counties. Find the sample variance.

4.2    3.6 2.8    3.0    3.4    2.9 5.6 2.9 1.7 2.5
Answer #1

Solution :

n = 10

\bar x = \sum x / n = 32.6 / 10 = 3.26

s2 =  \sum(x - \bar x)2 / n - 1  

= (4.2 - 3.26)2 + (3.6 - 3.26)2 + (2.8 - 3.26)2​​​​​​​ + (3 - 3.26)2​​​​​​​ + (3.4 - 3.26)2​​​​​​​ +

(2.9 - 3.26)2​​​​​​​ + (5.6 - 3.26)2​​​​​​​ + (2.9 - 3.26)2​​​​​​​ + (1.7 - 3.26)2​​​​​​​ + (2.5 - 3.26)2​​​​​​​ / 10 - 1

= 10.044 / 9

= 1.116

Sample variance = 1.116